# Cyber Attacks Detection, Prevention, and Source Localization in Digital Substation Communication using Hybrid Statistical-Deep Learning

### v1

@@ IV-E Attacker Considerations
From an attacker’s perspective, to deceive the proposed method several non-trivial challenges need to be overcome. First, estimating $F_{as}$ for a specific SV subscriber from a different device requires the estimation of the overall communication latency between the legitimate publisher and the targeted subscriber, and between the compromised device and the targeted subscriber. Furthermore, this estimation needs to keep into account also the timing synchronization errors. Second, to inject the SV frames at the exact time instant required to match the expected $F_{as}$, the compromised device needs to be executing a real-time operating system and be time synchronized with Precision Time Protocol (PTP). Moreover, to carry out the attack, a malicious script capable of publishing SV frames must be delivered and deployed on the compromised device. This deployment must occur stealthily, avoiding detection by both host-based and network-based IDSs; also, the script execution should not exhaust compromised device resources. Finally, to cause protection schemes to react to a fault, it is not sufficient to have only a small number of false measurements to be accepted. For instance, when overcurrent protection is implemented, the measured current needs to be above a certain threshold for at least 200 ms. This means that an attacker needs at least $FS/5$ consecutive malicious frames to be accepted to trigger the protection scheme. Thus, the injected malicious frames need to be consistently accepted for a non-negligible amount of time. All these required conditions combined make it extremely difficult for an attacker to successfully deceive the proposed method even in ideal conditions.
V Experimental Results and Discussion
@@
The proposed method was extensively tested on data acquired from three different IEC 61850-compliant testbeds. The first testbed consists of Real-Time Digital Simulator (RTDS) simulating an IEEE 5-bus system, a GTNETx2 card acting as SV publisher, a network switch, and three vIEDs connected in an HIL configuration and acting as SV subscribers. The second testbed is deployed by Elia transmission system operator in Belgium and consists of industrial-grade devices deployed in a Parallel Redundancy Protocol (PRP) configuration. In this second case, the deployed devices include an MU publishing an SV stream, two network switches, and an IED. Finally, the third testbed consists of a fully virtualized setup in which the digital substation communication network was emulated in Mininet. The setup consists of two HSR rings corresponding to the process and station buses, connected through a QuadBox, and containing 4 vMUs and 5 vIEDs, respectively. In the first two testbeds, the tests consisted in the execution of SV injection attacks from a compromised device connected to the network switch taking care of forwarding SV streams from the publisher to the subscriber. The injected false measurements represented an ongoing fault which successfully caused the (v)IEDs to react and issue an unintended command to open a circuit breaker.
The objective of the IPS was to prevent the malicious SV frames from being processed by the protection scheme while minimizing the number of legitimate frames being dropped. It should be noted that the IPS did not rely on the measurements content for frames classification, thus the difference between the legitimate and malicious measurements was not relevant for this performance evaluation and can be assumed to be zero. On the other hand, the third testbed was used to monitor network traffic during MitM attacks; this data was then used for the training and testing of the attack source localization module. To acquire the data needed for the validation of the proposed method, Wireshark was run on each (v)IED belonging to the three testbeds. Fig. 3 shows how EMG distribution fits the measured $F_{as}$ values at one vIED deployed in the third testbed. As can be seen, the EMG distribution provides better fitting capabilities compared to Exponential and Gaussian distributions.
@@
In Table III, the main information about the monitored SV streams in the first two testbeds during normal conditions is reported. This information includes the number of frames received per second ($FS$), measured $F_{as}$ mean and standard deviation, and the overall time synchronization error due to the lack of time synchronization with the PTP master clock. It should be noted that to comply with IEC 61850-9-3, each IED should guarantee a maximum holdover of ± 0.2 $\mu s/s$, thus implying a maximum time synchronization divergence among two devices of ± 0.4 $\mu s/s$. Given the holdovers experienced during our tests, it is fair to claim that the proposed method was tested in the most critical conditions that could be found in a real digital substation scenario. In the following, each component performance is evaluated in detail.
Figure 3: Demonstration of EMG distribution fitting capabilities.
@@
$FS$ [$frames/s$] 4800 4800 4800 4000
$\mathbb{E}[F_{as}]$ [$ms$] -7.08 -4.60 0.447 -113.30
$F_{as}$ std. dev [$\mu s$] 53 123 104 25
Holdover [$\mu s$] -1.399 1.658 -0.032 -0.371
V-A Intrusion Prevention Evaluation
To effectively evaluate the performance of the proposed intrusion prevention method, the TPR, FPR, and F1-score metrics are considered. In fact, for a fair evaluation it is crucial to keep into consideration and maximize the TPR while minimizing the FPR. This is because a considerable number of consecutive false negatives could lead the (v)IED to issue unintended control commands to circuit breakers. Conversely, if many legitimate SV frames are discarded because they are wrongly classified, the protection algorithm is prevented from operating properly and timely. Before testing the IPS against SV injection attacks, its performances in terms of false positives during normal operating conditions are evaluated. During normal conditions, the IPS caused FPR of 0.07%, 0.35%, 0.32%, and 0.005% on vIED1, vIED2, vIED3, and IED1, respectively. These false positives were mainly caused by frames being received with an unexpected latency greater than 3 ms. Indeed, as can be noted in Table III, higher FPR resulted in (v)IEDs experiencing higher levels of communication jitter, i.e., higher variance of $F_{as}$. Given that the IEC 61850 standard mandates a maximum latency of 3 ms, these frames should have been discarded anyway by the protection scheme and thus should not be considered an actual misbehavior of the proposed IPS. Subsequently, malicious SV streams were injected into the network from a compromised machine connected to the network switch in the first and second testbeds. Due to the attacker’s challenges discussed in Section IV, perfectly matching the expected $F_{as}$ was not feasible. In fact, after several trials and adjustments, the minimum difference between the frame arrival time shift expected by the SV subscriber ($F_{as}^{leg}$) and frame arrival time shift of the frames injected by the attacker ($F_{as}^{mal}$) was of -178 $\mu s$. The difference between $F_{as}^{leg}$ and $F_{as}^{mal}$, which will be referred to as $F_{as}^{(mal-leg)}$, can be seen as the attacker injection time shift error. As can be appreciated in Table IV, even with such small attacker injection time shift errors, the IPS provided FPRs lower than 0.77% and TPRs higher than 98.5%.
TABLE IV: IPS Performances Against SV Injection Attacks Performed on the Two Testbeds.
@@ V-A Intrusion Prevention Evaluation
To accurately estimate the IPS performance in function of $F_{as}^{(mal-leg)}$, a synthetic dataset was generated starting from the original one. In the synthetic dataset, the original legitimate frames were duplicated, and then the duplicated frames $F_{as}$ was varied such that the values of $F_{as}^{(mal-leg)}$ were included within ± 600 $\mu s$. These duplicated frames represented the malicious frames injected by the attacker. As can be observed in Fig. 4, the IPS provided great prevention capabilities down to an attacker injection error of less than ± 100 $\mu s$, with FPRs lower than 1% and F1-scores above 95%. Higher $F_{as}^{(mal-leg)}$ absolute values were not considered because they would lead to F1-scores close to 100%. As expected, the limitation of the proposed IPS comes into play when the attacker matches the expected legitimate $F_{as}$. In that case, the information advantage is lost, and legitimate and malicious frames have a 50% probability of being accepted or discarded. Final consideration regards the IPS added latency and throughput performances. Throughout the tests, it was verified that the IPS introduced an additional latency of less than 100 $\mu s$ and sustained a throughput exceeding 100,000 frames per second. These performances satisfy the stringent timing requirements specified in IEC 61850-5 and support the highest SV frames publishing rate of 96,000 frames per second, as defined in IEC 61869-9.
Figure 4: Intrusion prevention method FPR and F1-score with varying attacker injection error ($F_{as}^{(mal-leg)}$) on the four (v)IEDs.
V-B MitM Detection Evaluation
The MitM attacks detection component was evaluated on four synthetic datasets generated starting from the ones acquired from the first two testbeds. It was assumed that the attacker, by compromising a forwarding device and performing a MitM attack against the subscriber, could drop the legitimate SV stream and inject the malicious one. Thus, the SV subscriber received exactly the expected number of SV frames, i.e., $FS$, but, due to the MitM attack, the frames arrival time statistical properties were affected. This assumption was then validated on the fully virtualized setup, where the MitM attacks were practically performed. To generate the synthetic anomalous acquisitions, half of the legitimate data was modified by adding random variations in the measured $F_{as}$ mean, standard deviation, and skewness. These random variations are referred to as $\Delta m$, $\Delta s$, and $\Delta\gamma_{1}$, with values drawn from Uniform distributions with lower and upper bounds of ± 300 $\mu s$, ± 10%, and ± 5%, respectively. After selecting a detection threshold guaranteeing FPRs lower than 0.15% and 0.01% in the first and second testbeds, respectively, the MitM component provided TPRs of 95.15%, 95.99%, 93.85%, and 97.66%, for vIED1, vIED2, vIED3, and IED1, respectively. Still, it must be kept in consideration that in the generated synthetic dataset, some malicious samples were indistinguishable from the legitimate ones, due to random choosing of $\Delta m$, $\Delta s$, and $\Delta\gamma_{1}$ values close to zero. In Fig. 5, the evaluation results for the second testbed IED with varying $\Delta m$ and $\Delta s$ are depicted. Similar consideration can be drawn for the first testbed vIEDs. As can be appreciated, the MitM detection method provided an F1-score higher than 99.9% down to an $\Delta m$ variation of 20 $\mu s$. This means that, to remain undetected, an attacker needs to alter the $F_{as}$ measured by less than 20 $\mu s$ on average. Moreover, even if a lower alteration is caused on the expected $F_{as}$ mean, variations in the measured standard deviation or skewness of $F_{as}$ still provide an information advantage to detect the ongoing MitM attack.
@@ V-B MitM Detection Evaluation
TABLE V: Separation of Scenarios Among Training and Testing Dataset.
@@ V-C Attack Source Localization Evaluation
This last component was validated on network traffic acquired from the third testbed. To perform MitM attacks, each vIED in the HSR ring was compromised one at a time, its NICs’ HSR bridge was disabled, and a C script was used to monitor, tamper, and forward SV frames. The network traffic was monitored with Wireshark at each vIED NIC. The network traffic was acquired during four different experimental scenarios. Each experimental scenario consisted of one case representing normal operating conditions and five cases with a MitM attack in place in one vIED, thus leading to 24 acquisitions in total to be used for the training and validation of the DL model. As for the validation of the MitM detection component, measured $F_{as}$ mean, standard deviation, and skewness were computed for each vIED, and then delivered to the attack source localization component. To further increase the realism of the experiments, the compromised vIED was allowed to report either normal or abnormal statistical metrics from both NICs. The DL model was trained on normal and anomalous cases of some experimental scenarios and then tested on normal and anomalous cases of other scenarios as reported in Table V; this allowed to further prove the generalization capabilities of the proposed DL-based method. In fact, in a real deployment scenario, only normal operating conditions can usually be acquired from the field and be used for training. However, as demonstrated in our experiments, the training set can be expanded with anomalous data acquired from an emulated network. The resulting trained model is then effective in detecting and localizing unseen MitM attacks in the real OT communication network. As reported in Table VI, both in the training and testing datasets, the attack source localization method provided accuracy, precision, recall, and F1-scores higher than 99%. Furthermore, as can be appreciated in Fig. 6, the number of false positives was limited to 6 out of 2366 normal operating conditions samples, resulting in a FPR of 0.25%. Also, when the source of the attack was mislocalized, most of the time the error was limited to devices adjacent to the malicious one. This limits the number of devices that require further investigation after a cyber security incident.
@@
Figure 6: Attack source localization confusion matrix on the testing set. Label ”0” corresponds to normal conditions; labels from ”1” to ”5” denote which vIED is performing the MitM attack.
VI Conclusions

### v2

@@ III-E Attacker Considerations
From an attacker’s perspective, several challenges need to be overcome to deceive the proposed intrusion detection and prevention method. First, estimating $F_{as}$ for a specific SV subscriber from a different device in the network requires the estimation of the overall communication latency between the legitimate publisher and the targeted subscriber, and between the compromised device and the targeted subscriber. Furthermore, this estimation needs account for timing synchronization errors. Second, to inject the SV frames at the exact time instant required to match the expected $F_{as}$, the compromised device needs to be executing a real-time operating system and be time synchronized with Precision Time Protocol (PTP). Moreover, to carry out the attack, a malicious script capable of publishing SV frames must be delivered, installed, and executed with high privileges on the compromised device. This deployment must occur stealthily, avoiding detection by both host-based and network-based IDSs; also, the script execution should not exhaust compromised device resources. Finally, to cause protection schemes to react to a fault, it is not sufficient to have only a small number of false measurements to be accepted. For instance, when overcurrent protection is implemented, the measured current needs to be above a certain threshold for at least 200 ms. This means that an attacker needs at least $FS/5$ consecutive malicious frames to be accepted to trigger the protection scheme. Thus, the injected malicious frames need to be consistently accepted for a non-negligible amount of time. The combination of these conditions imposes considerable complexity on potential attacks, effectively preventing successful deception of the proposed method, even under ideal scenarios. In the remainder of the paper, it is assumed that the attacker has gained privileged access to a machine connected to the network switch forwarding the legitimate SV stream, or to a forwarding device itself when considering MitM attacks, and can execute arbitrary scripts with high privileges to directly access the compromised machine’s NIC and inject malicious SV frames.
IV Experimental Results and Discussion
(a) Three vIED in a HIL configuration and subscribing to the SV stream published by RTDS through the GTNETx2 card.
(b) Industrial MU, IED, and network switches in PRP configuration.
(c) Fully virtualized setup with vMUs, vIEDs, and QuadBox deployed in Mininet; the third vIED is shown as a compromised device for illustration purposes.
Figure 3: Schematic overview of the three experimental testbeds used to validate the proposed method.
@@ IV Experimental Results and Discussion
The proposed method was extensively tested on data acquired from three different IEC 61850-compliant testbeds. The first testbed consists of a Real-Time Digital Simulator (RTDS) simulating an IEEE 5-bus system, a GTNETx2 card acting as an SV publisher and GOOSE subscriber, a network switch, and three vIEDs connected to the network switch in a hardware-in-the-loop configuration, acting as SV subscribers. The second testbed is deployed by Elia transmission system operator in Belgium and consists of industrial-grade devices deployed in a Parallel Redundancy Protocol (PRP) configuration. In this second case, the deployed devices include an industrial MU publishing an SV stream, two IEC 61850-compliant network switches, and an IED. Finally, the third testbed consists of a fully virtualized setup in which the digital substation communication network was emulated in Mininet. The setup consists of two HSR rings, connected through a QuadBox, and containing 4 vMUs and 5 vIEDs, respectively. Fig. 3 depicts the architecture of the three testbeds.
The experiments in the first testbed involved both SV injection attacks from a compromised device connected to the network switch and MitM attacks against one vIED. The second testbed focused exclusively on SV injection attacks from a compromised device connected to one of the network switches. Finally, the third testbed was dedicated to performing MitM attacks from each vIED in the HSR ring, one at a time. During SV injection attacks, the attacker injected malicious measurements while spoofing the identity of the legitimate SV publisher, emulating persistent false fault conditions that triggered the (v)IEDs to incorrectly issue trip commands for opening circuit breakers. The IPS aimed to block these malicious SV frames from reaching the protection scheme while minimizing the drop of legitimate traffic. MitM attacks against the vIEDs verified that changing NIC configuration and forwarding and tampering SV frames in user-space through a highly optimized C script would still cause detectable changes in the statistical distribution of frame arrival times received by the targeted IED. Finally, data collected in the third testbed was used for training and testing the attack source localization module. In all cases, network traffic acquisitions were obtained by running Wireshark on each (v)IED and network switch across the three testbeds. Notably, the integrated method classified frames, detected MitM attack, and localized the compromised devices without inspecting measurement data, rendering differences between legitimate and malicious measurements irrelevant for this evaluation.
TABLE II: AIC, BIC, and KS test results for Gaussian, exponential, and EMG distributions fitted on SV streams on one (v)IED per testbed.
Testbed PDF AIC BIC $\times 10^{3}$] KS
vIED1 Guassian Exponential EMG 11.35 10.46 9.41 454 418 377 0.271 0.328 0.083
IED1 Guassian Exponential EMG 9.15 9.19 8.22 366 368 329 0.271 0.370 0.167
vIEDHSR Guassian Exponential EMG 12.03 10.77 10.61 24 22 22 0.223 0.103 0.082
Figure 4: Comparison of EMG, Gaussian, and exponential distribution fits to the measured $F_{as}$.
IV-A Frame Arrival Time Characterization
Fig. 4 illustrates how the EMG distribution aligns with the measured $F_{as}$ values at one vIED deployed in the third testbed. Also, Table II reports the Akaike’s Information Criterion (AIC) normalized by sample size, the Bayesian Information Criterion (BIC), and the Kolmogorov-Smirnov (KS) test results computed on the monitored SV streams for one (v)IED for each testbed. As observed, the EMG distribution achieves a better fit than Gaussian and exponential distributions across all testbeds. It is acknowledged that an exhaustive search over alternative PDFs could identify models with improved fit for this particular dataset. Nevertheless, such an approach would risk overfitting and undermine the physical interpretability of the modeled data. Therefore, the EMG distribution was selected based on domain knowledge, balancing statistical fit with interpretability and relevance to substation communication patterns, rather than adopting a brute-force optimization strategy.
@@ IV-A Frame Arrival Time Characterization
The main information about the monitored SV streams in the first two testbeds during normal conditions is reported in Table III. This information includes the number of frames received per second ($FS$), measured $F_{as}$ mean and standard deviation, and the overall time synchronization error due to the lack of time synchronization with the PTP master clock. It should be noted that to comply with IEC 61850-9-3, each IED should guarantee a maximum holdover of ± 0.2 $\mu s/s$, thus implying a maximum time synchronization divergence among two devices of ± 0.4 $\mu s/s$. Given the holdovers observed during our tests, it is reasonable to assert that the proposed method was evaluated under the most challenging conditions likely to be encountered in real digital substation environments.
IV-B Intrusion Prevention Evaluation
To effectively evaluate the performance of the proposed intrusion prevention method, the True Positive Rate (TPR), FPR, and F1-score metrics are considered. For a fair assessment, it is essential to keep into consideration and maximize the TPR while minimizing the FPR. This is because a considerable number of consecutive false negatives can lead the (v)IED to issue unintended control commands to circuit breakers. Conversely, if many legitimate SV frames are discarded because wrongly classified, the protection algorithm is prevented from operating properly and in a timely manner.
@@ IV-B Intrusion Prevention Evaluation
FSFS [$frames/s$] 4800 4800 4800 4000
$\mathbb{E}[F_{as}]$ [msms] -7.08 -4.60 0.447 -113.30
$F_{as}$ std. dev $\mu s$] 53 123 104 25
Holdover $\mu s/s$] -1.399 1.658 -0.032 -0.371
Before testing the IPS against SV injection attacks, its performance in terms of false positives during normal operating conditions is evaluated. During normal conditions, the IPS caused FPR of 0.07%, 0.35%, 0.32%, and 0.005% on vIED1, vIED2, vIED3, and IED1, respectively. These false positives were mainly caused by frames being received with an unexpected latency greater than 3 ms. Indeed, as can be noted in Table III, higher FPR resulted in (v)IEDs experiencing higher levels of communication jitter, i.e., higher variance of $F_{as}$. These frames exceed IEC 61850 latency limit of 3 ms and would be discarded by protection schemes, aligning with expected IPS behavior.
Subsequently, malicious SV streams were injected into the OT communication network from a compromised machine connected to the network switch in the first and second testbeds. Due to the attacker’s challenges discussed in Section III-E, perfectly matching the expected $F_{as}$ was not feasible. In fact, after several trials and adjustments, the minimum difference between the frame arrival time shift expected by the SV subscriber ($F_{as}^{leg}$) and frame arrival time shift of the frames injected by the attacker ($F_{as}^{mal}$) was -178 $\mu s$. The difference between $F_{as}^{leg}$ and $F_{as}^{mal}$, which will be referred to as $F_{as}^{(mal-leg)}$, can be seen as the attacker’s injection time shift error. As can be appreciated in Table IV, even with such small attacker injection time shift errors, the IPS provided FPRs lower than 0.77% and TPRs higher than 98.5%.
TABLE IV: IPS performances against SV injection attacks performed on the first two testbeds.
@@
Figure 5: FPR and F1-score of the intrusion prevention method as a function of attacker injection error on the four (v)IEDs deployed across the first and second testbeds.
To accurately estimate the IPS performance as a function of $F_{as}^{(mal-leg)}$, a synthetic dataset was generated from the original one. In this synthetic dataset, the original legitimate frames were duplicated, and the $F_{as}$ values of the duplicated frames were varied such that $F_{as}^{(mal-leg)}$ lies within ± 600 $\mu s$. These duplicated frames represent the malicious frames injected by the attacker. This approach allows to accurately identify method’s prevention boundaries by analyzing its performance in the relevant search space that reflects attacker capabilities.
As can be observed in Fig. 5, the IPS provided great prevention capabilities down to an attacker injection error of less than ± 100 $\mu s$, with FPRs lower than 1% and F1-scores above 95%. Higher $F_{as}^{(mal-leg)}$ absolute values were not considered because they would lead to F1-scores close to 100%. As expected, the limitation of the proposed IPS comes into play when the attacker matches the expected legitimate $F_{as}$. In that case, the information advantage is lost, and legitimate and malicious frames have a 50% probability of being accepted or discarded. Final consideration regards the IPS added latency and throughput performances. Throughout the tests, it was verified that the IPS introduced an additional latency of around 150 $\mu s$ and sustained a throughput exceeding 100,000 frames per second. These performances satisfy the stringent timing requirements specified in IEC 61850-5 and support the highest SV frames publishing rate of 96,000 frames per second, as defined in IEC 61869-9.
IV-C MitM Detection Evaluation
Experiments on the first testbed revealed that MitM attacks against a (v)IED produced significant changes in SV frame arrival-time statistics. Specifically, frame interception and tampering increased latency by 40.55 $\mu s$, jitter by 8.48%, and PDF skewness by 115.79%, on average. These results demonstrate that SV frame arrival-time statistics provide a reliable indicator for MitM attack detection.
@@ IV-C MitM Detection Evaluation
Consequently, the MitM attack detection component was evaluated on four synthetic datasets generated from the network traffic acquisitions from the first two testbeds, enabling the identification of the component’s detection boundaries. The attacker, by compromising a forwarding device and performing a MitM attack against the subscriber, could drop the legitimate SV stream and inject a malicious one in real-time. As a result, the SV subscriber received exactly the expected number of SV frames, i.e., $FS$, but the statistical properties of their arrival times were altered due to the MitM attack. This expected behaviour was experimentally verified in the third testbed, where the MitM attacks were practically performed.
To generate the synthetic anomalous acquisitions, half of the legitimate data was modified by introducing variations in the mean, standard deviation, and skewness of the measured $F_{as}$. These variations are referred to as $\Delta m$, $\Delta s$, and $\Delta\gamma_{1}$, with values drawn from uniform distributions with lower and upper bounds of ± 300 $\mu s$, ± 10%, and ± 5%, respectively. After selecting a detection threshold providing FPRs lower than 0.15% and 0.01% in the first and second testbeds, respectively, the MitM component provided TPRs of 95.15%, 95.99%, 93.85%, and 97.66%, for vIED1, vIED2, vIED3, and IED1, respectively. Still, it must be kept in consideration that in the generated synthetic dataset, some malicious samples were indistinguishable from the legitimate ones, due to random choosing of $\Delta m$, $\Delta s$, and $\Delta\gamma_{1}$ values close to zero. In Fig. 6, the evaluation results for the second testbed IED with varying $\Delta m$ and $\Delta s$ are depicted. Similar consideration can be drawn for the first testbed vIEDs. As can be appreciated, the MitM detection method provided an F1-score higher than 99.9% down to an $\Delta m$ variation of 20 $\mu s$. This means that, to remain undetected, an attacker needs to alter the $F_{as}$ measured by less than 20 $\mu s$ on average. Moreover, even if a lower alteration is caused on the expected $F_{as}$ mean, variations in the measured standard deviation or skewness of $F_{as}$ still provide an information advantage to detect the ongoing MitM attack.
IV-D Attack Source Localization Evaluation
TABLE V: Separation of scenarios among training and testing datasets.
@@ IV-D Attack Source Localization Evaluation
This last component was validated on network traffic acquired from the third testbed. To perform MitM attacks, each vIED in the HSR ring was compromised one at a time, its NICs’ HSR bridge was disabled, and a C script was used to monitor, tamper, and forward SV frames. The network traffic was monitored and acquired with Wireshark at each vIED NIC during four experimental scenarios. Each experimental scenario consisted of one case representing normal operating conditions and five cases with a MitM attack in place in one vIED, thus leading to 24 acquisitions in total to be used for the training and testing of the DL model. As for the validation of the MitM detection component, measured $F_{as}$ mean, standard deviation, and skewness were computed for each vIED, and then delivered to the attack source localization component. To further increase the realism of the experiments, the compromised vIED was allowed to report either normal or abnormal statistical metrics from both NICs. The DL model was trained on normal and anomalous cases from a subset of the experimental scenarios and then tested on normal and anomalous cases from different scenarios, as reported in Table V. This setup further demonstrates the generalization capabilities of the proposed DL-based method.
In fact, in a real deployment scenario, only normal operating conditions can usually be acquired from the field and used for training. However, as demonstrated in our experiments, the training set can be expanded with anomalous data acquired from an emulated network. The resulting trained model is then effective in detecting and localizing unseen MitM attacks in the real OT communication network. As reported in Table VI, the attack source localization method achieved accuracy, precision, recall, and F1-scores higher than 99% on both the training and testing datasets. Furthermore, as can be appreciated in Fig. 7, a FPR of 0.25% was obtained on the testing dataset. When the attack source was mislocalized, the error was most often confined to devices adjacent to the malicious one. This limits the number of devices that require further investigation after a cyber security incident.
@@
Figure 7: Confusion matrix for attack source localization component on the testing set. Label “0” corresponds to normal conditions; labels “1” to “5” indicate which vIED is performing the MitM attack in the third testbed.
IV-E Final Discussion and Considerations
Table VII summarizes the protection capabilities of different methods against the cyber attacks introduced in Section II-B. As can be appreciated, the method proposed in this paper addresses the limitations of conventional IDSs by effectively detecting and preventing more advanced attacks.
During flooding attacks, the proposed method discerns the legitimate and malicious frames and ensures that at most $FS$ frames per second are received by the protected IED, thus drastically reducing the targeted device’s computational load and successfully mitigating the attack. While conventional IDSs can detect SV frames reception rate anomalies, they lack mechanisms to forward only legitimate frames to IEDs.
For replay attack injecting previously sniffed SV frames, the arrival PDF and replay protection module detects and discards duplicated smpCnt values. Although alternative IDSs monitoring smpCnt can identify such attacks, they cannot differentiate legitimate frames from malicious duplicates. Further, IDSs relying on current and voltage measurements monitoring remain ineffective because injected values may remain consistent with normal power system operation.
As shown in the quantitative analysis of the experimental scenarios, spoofing attacks are prevented by discarding the SV frames with the lowest legitimacy likelihood. The inherent difficulty of accurately and consistently estimating $F_{as}$ prevents attackers from successfully mimicking legitimate traffic, thereby enabling effective detection and prevention of spoofing attacks. In contrast, other IDSs that rely on frame-rate monitoring can detect spoofing when both legitimate and malicious SV streams are received, but they cannot distinguish which frames are legitimate.
TABLE VII: Existing and proposed solutions’ protection capability against different types of attacks, where ✓ denotes effective protection, ✗ ineffective protection, and $\sim$ partial protection.
(1) (2) (3) (4) (5) (6)
Ustun et al. [27] $\sim$
Mo et al. [20] $\sim$
El Hariri et al. [7] $\sim$
Hussain et al. [16] $\sim$ $\sim$
Hong et al. [14] $\sim$
Delhomme et al. [6] $\sim$
Wannous et al. [28]
Hong et al. [13, 12] $\sim$
Eynawi et al. [9]
Manzoor et al. [19]
Our Solution
High smpCnt attacks, which trick replay protection by injecting frames with artificially elevated smpCnt field values, are similarly thwarted because the frame arrival time constraints remain unsatisfied, causing malicious frames to be discarded.
During MitM attacks, where legitimate frames are dropped and malicious ones are injected at exactly $FS$ frames per second, the arrival PDF and replay protection modules may initially accept malicious frames. However, the MitM detection module identifies the attack because matching the statistical properties, i.e., mean, variance, and skewness, of the legitimate SV stream presents substantial difficulty, particularly under realistic operating conditions. Prior IDS approaches cannot detect such attacks when the frame rate, smpCnt field, SV stream identifier, and inter-frame timing are all properly emulated.
Finally, when SV frame authentication is enabled, but the pre-shared authentication key has been compromised, the proposed method’s arrival PDF and replay protection component detects and prevents attacks failing to satisfy the expected frame arrival time shift constraint, ensuring similar performances as during spoofing attack scenarios.
V Conclusions